# Fast and accurate parameter estimation of high-redshift sources with the Einstein Telescope

### v1

@@ IV.2 Higher-order modes and precessing spins
where $Y_{-2}^{l,m}$ are the spin-weighted spherical harmonics of weight $-2$ (Blanchet, 2014; Mills and Fairhurst, 2021), which capture the angular dependence of the radiation, and $\vec{\lambda}$ encompasses the intrinsic parameters, such as component masses and spins. The terms $h_{\ell,m}$ are the GW modes (or multipoles), with the dominant quadrupole mode corresponding to $\ell=2$, and $m=\pm 2$. Higher-order modes (HOMs) appear for $\ell\geq 2$ and $|m|\neq 2$, and are also referred to as higher multipoles (Khan et al., 2020), higher spherical harmonics (Mills and Fairhurst, 2021), non-quadrupole or subdominant modes.
@@
In this work, we assume aligned spins, meaning the spins of the individual black holes are parallel to the orbital angular momentum of the binary system (Eq. 4), so no spin precession occurs (Apostolatos et al., 1994). In this case, the waveform can be written in the frequency domain using the stationary-phase approximation, where the HOM amplitudes for both polarizations depend on the inclination angle (Mills and Fairhurst, 2021).
@@
When more than one mode is observed, the measurement of the inclination angle can be improved (Abbott et al., 2020a; Kalaghatgi et al., 2020), reducing the correlation with the luminosity distance. In addition, detecting HOMs leads to a more accurate measurement of the properties of the source, in general, as they help to model the GW signal more accurately (Kalaghatgi et al., 2020; Abbott et al., 2020a, b; Marsat et al., 2021; Baral et al., 2025; Yi et al., 2025; Pitte et al., 2023). HOMs can aid in early warning detections since they have longer in-band signal duration in the time domain with respect to the dominant mode (Kapadia et al., 2020). Lastly, they can be used to test general relativity (Puecher et al., 2022).
@@
The contribution of HOMs relative to the leading mode depends largely on the system asymmetries. Their amplitude is higher in binaries with large mass ratios, making them particularly relevant for neutron star–black hole binaries (Vitale and Chen, 2018; Zhang et al., 2023), in binaries with large misaligned spins (Singh et al., 2024), and in sources viewed at large inclination angles, close to edge-on ($\theta_{\mathrm{jn}}\sim\pi/2$; Mills and Fairhurst, 2021).
@@
If spin precession is allowed, the symmetry $h_{\ell,-m}=(-1)^{\ell}h^{*}_{\ell,m}$ is broken (Boyle et al., 2014), meaning the negative $m$-modes are no longer directly related to the positive $m$-modes. Precession also causes modulations of the amplitude and frequency of the plus and cross polarizations (Calderón Bustillo et al., 2017), leading to an effective mode mixing (Schmidt et al., 2012; Ossokine et al., 2013). This allows for an even more accurate measurement of the inclination angle, helping to resolve the degeneracy with the luminosity distance, which in turn significantly improves distance and sky localization estimates when both HOMs and precessing spins are detected (London et al., 2018; Chen et al., 2019; Vitale and Chen, 2018).
@@
Including all spin parameters in our analysis, therefore allowing for spin precession, is crucial for better constraining source parameters and resolving degeneracies. However, spin precession introduces frequency-domain modulations, making the data more complex. To address this, we need to improve our method performance, which may require training Dingo-IS with group-equivariant NPE (GNPE; Dax et al., 2021) or flow matching posterior estimation (FMPE; Wildberger et al., 2023).
@@
GNPE exploits physical symmetries in GW signals, such as arrival time differences between detectors, by using an additional neural network to learn and apply time shifts to both parameters and data. This enables Dingo-IS to focus on smaller time shifts, which correspond to minor phase shifts in the frequency domain, simplifying the data stream and improving inference. Flow matching, a recent generative modeling technique, enhances the flexibility of network architectures for handling complex data (Lipman et al., 2023).
@@
We generate waveforms using the IMRPhenomXPHM approximant (Pratten et al., 2021), including all available HOMs and, separately, only the dominant quadrupole mode $(\ell,m)=[(2,2),(2,-2)]$. We perform inference with the same sampling setup, using $5000$ live points on four CPUs. The sampling time differs significantly: about 10 hours for HOMs with $3.7\times 10^{7}$ likelihood evaluations, while the dominant mode requires 13 days and 18 hours with $8.5\times 10^{8}$ likelihood evaluations. This is because the quadrupole mode cannot resolve degeneracies, making sampling slower and PE less precise, as shown in Fig. 11. These results align with previous findings for current gravitational wave interferometers (Abbott et al., 2020a) and LISA (Marsat et al., 2021), confirming the importance of HOMs for PE with ET (Fairhurst et al., 2024).
@@
Figure 12: Marginalized one- and two-dimensional posterior distributions for luminosity distance ($d_{\mathrm{L}}$), azimuth ($\lambda$), altitude ($\beta$), and inclination angle ($\theta_{\mathrm{jn}}$) for Event 1 with zero noise. Waveform generation is done using all higher-order modes (HOMs) available with IMRPhenomXPHM approximant. The posterior distributions correspond to three cases: HOM A (dark green) with aligned spins, HOM P (yellow) where precessing spin parameters are sampled, and HOM Pinj (grey) where precessing spin parameters are fixed to their injected values. Contours indicate the 68% and 95% credible regions. Vertical and horizontal lines mark the true injected values. See Sec. IV.2 for details.
@@
We analyze a second injection, again without noise, using the same parameters of Event 1 but modifying the spin angles. In this case, we introduce precessing spins with the following parameters: $\theta_{1}=0.5$ rad, $\theta_{2}=1.0$ rad, $\phi_{12}=1.7$ rad, and $\phi_{jl}=0.3$ rad. Here, $\phi_{12}$ is the angle between the projections of the two spin vectors in the orbital plane, while $\phi_{jl}$ is the angle between the orbital and total angular momentum (see Fig. 7 of Dupletsa et al. 2025). We consider two inference setups: in the first, we fix the spin angles to their injected values by using delta-function priors; in the second, we allow $\theta_{1}$, $\theta_{2}$, $\phi_{12}$, and $\phi_{jl}$ to be freely sampled.
@@
The outcome of this test is shown in Fig. 12. When the precessing spin parameters are perfectly known, the degeneracy between the inclination angle and luminosity distance is resolved. As a consequence, the number of sky modes discussed in Sec. IV.1 is reduced to four. In particular, measuring sufficiently well spin precession removes the degeneracy between inclination angle ($\theta_{\mathrm{jn}}$) and altitude ($\beta$), which means the sky modes remain degenerate in the azimuth angle ($\lambda$). However, accurately constraining the precessing spin parameters is not guaranteed, even for an high optimal SNR signal ($\rho\sim 83$) like the one considered in our case. This aligns with previous research showing that the ability to estimate precession depends heavily on the specific characteristics of the source (Mills and Fairhurst, 2021; Rosenberg et al., 2024), including signal duration, with longer signals providing better constraints (Romero-Shaw et al., 2023).
IV.3 Astrophysical population
@@
We estimate the parameters of GW events extracted from an astrophysical population, which is compatible with the prior ranges listed in Tab. 1. Specifically, we focus on BBHs formed from Pop. III stars, which are believed to form and merge at high redshift (Liu et al., 2024; Mestichelli et al., 2024). Our analysis includes one year of BBH mergers predicted by the LOG1 model (Costa et al., 2023; Santoliquido et al., 2023). This model assumes a flat-in-log initial mass function (Stacy and Bromm, 2013; Tanikawa et al., 2021; Prole et al., 2022) and adopts the mass ratio, eccentricity and orbital period of zero-age main sequence progenitor stars from Sana et al. (2012).
Single and binary stellar evolution processes are modeled using the population synthesis code sevn (Iorio et al., 2023). The merger rate density of Pop. III BBHs is computed with $\mathcal{R}$ate (Santoliquido et al., 2020, 2021), which combines population synthesis results with the star formation rate density model of Pop. III stars from Hartwig et al. (2022). For further details, see Santoliquido et al. (2024).
@@
To create our injected BBHs from Pop. III stars, we take only the chirp mass, mass ratio, and luminosity distance predicted with the LOG1 model. The other waveform parameters are sampled from their prior distributions (Sec. II.2). The LOG1 model predicts about 690 BBH mergers per year, regardless of whether they are detectable with ET-$\Delta$. We processed all of them with Dingo-IS, except for 14 sources with an injected detector-frame chirp mass $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}<40$ M and 1 source with an injected luminosity distance $d^{\mathrm{inj}}_{\mathrm{L}}<5\times 10^{3}$ Mpc. For each source, we extracted $2\times 10^{4}$ samples. Fig. 13 confirms the results shown in Fig. 7, where injections where sampled from the priors distribution. The average and median sample efficiency are 11.4 % and 8.1%, respectively, with $\sim 75$% of sources with sample efficiency $>1\%$. In particular, sample efficiency is higher for sources merging at high redshift, reaching an average sample efficiency of 16% at $z>10$.
@@
Beyond BBHs formed from Pop. III stars, intermediate-mass black holes (IMBHs) represent another class of objects with masses in the range specified in Tab. 1 (Greene et al., 2020; Askar et al., 2023). In the $10^{2}$$10^{3}$ M range, IMBHs are thought to primarily form through stellar collisions (Portegies Zwart and McMillan, 2002; Giersz et al., 2015; Mapelli, 2016; Arca Sedda et al., 2021, 2023), hierarchical BH mergers (Miller and Hamilton, 2002; Gültekin et al., 2004; Antonini et al., 2019; Arca Sedda et al., 2021), close interactions between stars and BHs (Stone et al., 2017; Rizzuto et al., 2023; Arca Sedda et al., 2023), or a combination of these processes (Giersz et al., 2015; Arca Sedda et al., 2023). However, due to limited observational evidence, the mechanisms governing IMBH formation and growth remain unclear. In this context, ET and 3G detectors will provide crucial insights into the nature of IMBHs (Arca Sedda et al., 2020; Colpi et al., 2024; Ajith et al., 2025; Abac et al., 2025; Reali et al., 2024).
IV.4 Implications
@@
The results shown in this work have important implications. For instance, although ET may observe BBH mergers up to $z\sim 100$ (Ng et al., 2021, 2022a; Kalogera et al., 2021; Yi et al., 2022; Ng et al., 2022b; Branchesi et al., 2023), several studies have highlighted the difficulty ET will face in determining the luminosity distance distribution of high-redshift sources (Iacovelli et al., 2022; Mancarella et al., 2023; Marcoccia et al., 2024; Santoliquido et al., 2024). However, these studies predominantly used the FIM approximation. We expect that their conclusions could be more optimistic if PE were done with NPE, as presented in this work (see Fig. 8).
@@
NPE demonstrates the realistic sky localization performance of the triangular ET for short-lived high-redshift sources, showing that in the best-case scenario, the uncertainties are up to 8 times larger than those predicted using the FIM (see Fig. 8). This might have significant implications for dark siren cosmology, which involves statistically inferring the redshift from potential host galaxies within the estimated GW localization volume (Schutz, 1986; Del Pozzo, 2012; Chen et al., 2018; Libanore et al., 2021; Gair et al., 2023; Bosi et al., 2023; Borghi et al., 2024). Several studies have predicted the ability of ET to constrain at percent level cosmological parameters such as $H_{0}$ and $\Omega_{m}$ (Muttoni et al., 2022, 2023; Chen et al., 2024; Califano et al., 2025) primarily relying on sources at $z<1$, where galaxy catalogs are expected to be complete. These studies used the FIM to estimate the uncertainty in sky localization. Their results could be more optimistic than they would be with NPE, unless the sky localization is evaluated considering a network of 3G detectors.
@@
Sky modes will only be seen for short-lived sources, where antenna patterns are time-independent. For low-mass sources, such as binary neutron stars (BNSs), sky modes will be resolved once Earth’s rotation is taken into account during parameter estimation (Chan et al., 2018; Nitz and Dal Canton, 2021; Li et al., 2022, 2024; Singh and Bulik, 2022).
IV.5 Energy cost
@@
Any improvement in PE for 3G detectors, such as the NPE method presented here, must be both time-efficient to meet the scientific requirements of ET and energy-efficient to ensure sustainability (Wouters et al., 2024). This is important not only for making scientific research more environmentally friendly but also for understanding the energy demands of PE in preparation for future ET facilities, including power supply needs.
@@
Training takes 6 days using one GPU and 32 CPUs, consuming 970 kWh. Each injection is processed with 8 CPUs for about 1.7 minutes on average for a total of $\sim 29$ hours, consuming up to 45 kWh. In total, training and inference require 1015 kWh. For comparison, an injection processed with Bilby running on 4 CPUs, with nlive=2000, and using the nessai sampler takes about 5 hours on average, resulting in an estimated total energy consumption of 4000 kWh for all processed injections.
@@
IV.6 Caveats
A major drawback of likelihood-free inference methods like Dingo-IS is that if the training settings need to be changed, such as switching the waveform approximant, a new neural network must be trained using simulations generated with the updated settings. Training is not only the most time-consuming but also the most energy-intensive process of using NPE, as discussed in Sec. IV.5.
To accelerate training and improve convergence of Dingo-IS, we made a few simplifying assumptions. We set a low-frequency limit of $f_{\mathrm{min}}=6$ Hz, even though ET is expected to be sensitive down to $\sim 2$ Hz. Reducing $f_{\mathrm{min}}$ allows for processing longer signals—up to 64 s for mergers happening at $f<f_{\mathrm{max}}=256$ Hz—but requires a finer frequency resolution of $\mathrm{d}f=1/64$ Hz. This increases the input array size by a factor of 8, making training more challenging. Methods for processing longer signal with Dingo-IS have been developed in Dax et al. (2025), where they include the use of multibanding and heterodyning to handle BNS signals up to one hour in length.
Additionally, we did not consider contamination from overlapping signals or transient non-Gaussian noise, which could complicate inference (Raymond et al., 2024). Addressing these limitations in future work will help establish NPE as a robust method for processing the full range of detectable sources with ET.
A comprehensive comparison of full PE performance across different ET designs is planned for future work, including an investigation of the potential benefits of the 2L configuration, where the increased baseline between detectors may help reduce multimodalities in sky localization (Fairhurst, 2009, 2018; Grover et al., 2014; Berry et al., 2015).

### v2

@@ IV.3 Higher-order modes and precessing spins
where $Y_{-2}^{l,m}$ are the spin-weighted spherical harmonics of weight 2-2 (Blanchet:2013haa; Mills:2020thr), which capture the angular dependence of the radiation, and $\vec{\lambda}$ encompasses the intrinsic parameters, such as component masses and spins. The terms $h_{\ell,m}$ are the GW modes (or multipoles), with the dominant quadrupole mode corresponding to $\ell=2$, and $m=\pm 2$. Higher-order modes (HOMs) appear for $\ell\geq 2$ and $|m|\neq 2$, and are also referred to as higher multipoles (Khan:2019kot), higher spherical harmonics (Mills:2020thr), nonquadrupole or subdominant modes.
@@
In this work, we assume aligned spins, meaning the spins of the individual black holes are parallel to the orbital angular momentum of the binary system (Eq. 4), so no spin precession occurs (Apostolatos:1994mx). In this case, the waveform can be written in the frequency domain using the stationary-phase approximation, where the HOM amplitudes for both polarizations depend on the inclination angle (Mills:2020thr).
@@
When more than one mode is observed, the measurement of the inclination angle can be improved (LIGOScientific:2020stg; Kalaghatgi:2019log), reducing the correlation with the luminosity distance. In addition, detecting HOMs leads to a more accurate measurement of the properties of the source, in general, as they help to model the GW signal more accurately (Kalaghatgi:2019log; LIGOScientific:2020stg; LIGOScientific:2020zkf; Marsat:2020rtl; Baral:2025geo; Yi:2025pxe; Pitte:2023ltw). HOMs can aid in early warning detections since they have longer in-band signal duration in the time domain with respect to the dominant mode (Kapadia:2020kss). Lastly, they can be used to test general relativity (Puecher:2022sfm).
@@
The contribution of HOMs relative to the leading mode depends largely on the system asymmetries. Their amplitude is higher in binaries with large mass ratios, making them particularly relevant for neutron star–black hole binaries (Vitale:2018wlg; Zhang:2023ceh), in binaries with large misaligned spins (Singh:2023aqh), and in sources viewed at large inclination angles, close to edge-on ($\theta_{\mathrm{jn}}\sim\pi/2$; Mills:2020thr).
@@
If spin precession is allowed, the symmetry $h_{\ell,-m}=(-1)^{\ell}h^{*}_{\ell,m}$ is broken (Boyle:2014ioa), meaning the negative $m$-modes are no longer directly related to the positive $m$-modes. Precession also causes modulations of the amplitude and frequency of the plus and cross polarizations (CalderonBustillo:2016rlt), leading to an effective mode mixing (Schmidt:2012rh; Ossokine:2013zga). This allows for an even more accurate measurement of the inclination angle, helping to resolve the degeneracy with the luminosity distance, which in turn significantly improves distance and sky localization estimates when both HOMs and precessing spins are detected (London:2017bcn; Chen:2018omi; Vitale:2018wlg).
@@
Including all spin parameters in our analysis, thereby allowing for spin precession, is crucial for better constraining source parameters and resolving degeneracies. However, spin precession introduces frequency-domain modulations, making the data more complex. To address this, we need to improve our method performance, which may require training Dingo-IS with group-equivariant NPE (GNPE; Dax:2021myb) or flow-matching posterior estimation (FMPE; 2023mla..confE..34W).
@@
GNPE exploits physical symmetries in GW signals, such as arrival time differences between detectors, by using an additional neural network to learn and apply time shifts to both parameters and data. This enables Dingo-IS to focus on smaller time shifts, which correspond to minor phase shifts in the frequency domain, simplifying the data stream and improving inference. Flow matching, a recent generative modeling technique, enhances the flexibility of network architectures for handling complex data (lipman2023flowmatchinggenerativemodeling).
@@
We generate waveforms using the IMRPhenomXPHM approximant (PhysRevD.103.104056), including all available HOMs and, separately, only the dominant quadrupole mode $(\ell,m)=[(2,2),(2,-2)]$. We perform inference with the same sampling setup, using $5000$ live points on four CPUs. The sampling time differs significantly: about 10 hours for HOMs with $3.7\times 10^{7}$ likelihood evaluations, while the dominant mode requires 13 days and 18 hours with $8.5\times 10^{8}$ likelihood evaluations. This is because the quadrupole mode cannot resolve degeneracies, making sampling slower and PE less precise, as shown in Fig. 12. These results align with previous findings for current gravitational wave interferometers (LIGOScientific:2020stg) and LISA (Marsat:2020rtl), confirming the importance of HOMs for PE with ET (Fairhurst:2023beb).
@@
Figure 13: Marginalized one- and two-dimensional posterior distributions for luminosity distance ($d_{\mathrm{L}}$), azimuth ($\lambda$), altitude ($\beta$), and inclination angle ($\theta_{\mathrm{jn}}$) for Event 1 with zero noise. Waveform generation is done using all higher-order modes (HOMs) available with IMRPhenomXPHM approximant. The posterior distributions correspond to three cases: HOM A (dark green) with aligned spins, HOM P (yellow) where precessing spin parameters are sampled, and HOM Pinj (grey) where precessing spin parameters are fixed to their injected values. Contours indicate the 68% and 95% credible regions. Vertical and horizontal lines mark the true injected values. See Sec. IV.3 for details.
@@
We analyze a second injection, again without noise, using the same parameters of Event 1 but modifying the spin angles. In this case, we introduce precessing spins with the following parameters: $\theta_{1}=0.5$ rad, $\theta_{2}=1.0$ rad, $\phi_{12}=1.7$ rad, and $\phi_{jl}=0.3$ rad. Here, $\phi_{12}$ is the angle between the projections of the two spin vectors in the orbital plane, while $\phi_{jl}$ is the angle between the orbital and total angular momentum (see Fig. 7 of Dupletsa:2024gfl). We consider two inference setups: in the first, we fix the spin angles to their injected values by using delta-function priors; in the second, we allow $\theta_{1}$, $\theta_{2}$, $\phi_{12}$, and $\phi_{jl}$ to be freely sampled.
@@
The outcome of this test is shown in Fig. 13. When the precessing spin parameters are perfectly known, i.e., when the spin angles are fixed to their injected values, the degeneracy between the inclination angle and luminosity distance is resolved. As a consequence, the number of sky modes discussed in Sec. IV.1 is reduced to four. In particular, measuring sufficient spin precession removes the degeneracy between inclination angle ($\theta_{\mathrm{jn}}$) and altitude ($\beta$), which means the sky modes remain degenerate in the azimuth angle ($\lambda$). However, accurately constraining the precessing spin parameters is not guaranteed (see Fig. 13), even for an high optimal SNR signal ($\rho\sim 83$) like the one considered in our case. This aligns with previous research showing that the ability to estimate precession depends heavily on the specific characteristics of the source (Mills:2020thr; 2024AAS...24345609R), including signal duration, with longer signals providing better constraints (Romero-Shaw:2022fbf).
IV.4 Astrophysical population
Figure 14: Injections from an astrophysical distribution of binary black holes (BBHs) formed from Population III stars, shown as a function of detector-frame chirp mass ($x$-axis) and luminosity distance (left $y$-axis), with corresponding redshift values (right $y$-axis). Each point is color-coded by the sample efficiency, and its marginal distribution is shown as a gray histogram above. The gray lines represent constant chirp masses in the source frame, with values of $\mathcal{M}_{s}=5,\,10,\,15,\,25,\,40$ M. For more details, see Sec. IV.4.
We estimate the parameters of GW events extracted from an astrophysical population, which is compatible with the prior ranges listed in Table 1. Specifically, we focus on BBHs formed from Pop. III stars, which are believed to form and merge at high redshift (Liu:2024mkh; Mestichelli:2024djn). Our analysis includes one year of BBH mergers predicted by the LOG1 model (costa2023; santoliquido2023). This model assumes a flat-in-log initial mass function (Stacy:2012iz; Tanikawa:2020cca; Prole:2021nym) and adopts the mass ratio, eccentricity and orbital period of zero-age main sequence progenitor stars from Sana:2012px.
@@ IV.4 Astrophysical population
To create our injected BBHs from Pop. III stars, we take only the chirp mass, mass ratio, and luminosity distance predicted with the LOG1 model. The other waveform parameters are sampled from their prior distributions (Sec. II.2). The LOG1 model predicts about 690 BBH mergers per year, regardless of whether they are detectable with ET-$\Delta$. We processed all of them with Dingo-IS, except for 14 sources with an injected detector-frame chirp mass $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}<40$ M and one source with an injected luminosity distance $d^{\mathrm{inj}}_{\mathrm{L}}<5\times 10^{3}$ Mpc. For each source, we extracted $2\times 10^{4}$ samples. Figure 14 confirms the results shown in Fig. 7, where injections were sampled from the priors distribution. The average and median sample efficiency are 11.4 % and 8.1%, respectively, with $\sim 75$% of sources with sample efficiency $1\%$. In particular, sample efficiency is higher for sources merging at high redshift, reaching an average sample efficiency of 16% at $z>10$.
@@
Beyond BBHs formed from Pop. III stars, intermediate-mass black holes (IMBHs) represent another class of objects with masses in the range specified in Table 1 (2020ARA&A..58..257G; 2023arXiv231112118A). In the $10^{2}$$10^{3}$ M range, IMBHs are thought to primarily form through stellar collisions (2002ApJ...576..899P; 2015MNRAS.454.3150G; 2016MNRAS.459.3432M; 2021A&A...652A..54A; 2023MNRAS.526..429A), hierarchical BH mergers (2002MNRAS.330..232C; 2004ApJ...616..221G; 2019MNRAS.486.5008A; 2021A&A...652A..54A), close interactions between stars and BHs (2017MNRAS.467.4180S; 2023MNRAS.521.2930R; 2023MNRAS.526..429A), or a combination of these processes (2015MNRAS.454.3150G; 2023MNRAS.526..429A). However, due to limited observational evidence, the mechanisms governing IMBH formation and growth remain unclear. In this context, ET and 3G detectors will provide crucial insights into the nature of IMBHs (2020CQGra..37u5011A; 2024arXiv240207571C; 2025JCAP...01..108A; 2025arXiv250312263A; Reali:2024hqf).
IV.5 Implications
@@
The results shown in this work have important implications. For instance, although ET may observe BBH mergers up to $z\sim 100$ (Ng:2020qpk; Ng:2021sqn; Kalogera:2021bya; Yi:2022tzs; Ng:2022agi; Branchesi:2023mws), several studies have highlighted the difficulty ET will face in determining the luminosity distance distribution of high-redshift sources (Iacovelli:2022bbs; Mancarella:2023ehn; Marcoccia:2023khb; Santoliquido:2024oqs). However, these studies predominantly used the FIM approximation. We expect that their conclusions could be more optimistic if PE were done with NPE, as presented in this work (see Fig. 8).
@@
NPE demonstrates the realistic sky localization performance of the triangular ET for short-lived high-redshift sources, showing that for well-localized sources, where $\Delta\Omega_{90\%}<100$ deg2, the uncertainties are up to 8 times larger than those predicted using the FIM (see Fig. 8). This might have significant implications for dark siren cosmology, which involves statistically inferring the redshift from potential host galaxies within the estimated GW localization volume (Schutz:1986gp; DelPozzo:2011vcw; Chen:2017rfc; Libanore:2020fim; Gair:2022zsa; Bosi:2023amu; Borghi:2023opd). Several studies have predicted the ability of ET to constrain at percent level cosmological parameters such as $H_{0}$ and $\Omega_{m}$ (Muttoni:2021veo; Muttoni:2023prw; Chen:2024gdn; Califano:2025qbx) primarily relying on sources at $z<1$, where galaxy catalogs are expected to be complete. These studies used the FIM to estimate the uncertainty in sky localization. Their results could be more optimistic than they would be with NPE, unless the sky localization is evaluated considering a network of 3G detectors.
@@
Sky modes will only be seen for short-lived sources, where antenna patterns are time-independent. For low-mass sources, such as binary neutron stars (BNSs), which could last hours in the ET frequency band, we expect sky modes will be resolved once Earth’s rotation is taken into account during parameter estimation (Chan:2018csa; Nitz:2021pbr; Li:2021mbo; Li:2024ruk; Singh:2021bwn).
IV.6 Energy cost
@@
Any improvement in PE for 3G detectors, such as the NPE method presented here, must be both time-efficient to meet the scientific requirements of ET and energy-efficient to ensure sustainability (2024PhRvD.110h3033W). This is important not only for making scientific research more environmentally friendly but also for understanding the energy demands of PE in preparation for future ET facilities, including power supply needs.
@@
Training takes six days using one GPU and 32 CPUs, consuming 970 kWh. Each injection is processed with eight CPUs for about 1.7 minutes on average for a total of $\sim 29$ hours, consuming up to 45 kWh. In total, training and inference require 1015 kWh. For comparison, an injection processed with Bilby running on four CPUs, with nlive=2000, and using the nessai sampler takes about five hours on average, resulting in an estimated total energy consumption of 4000 kWh for all processed injections.
@@
IV.7 Caveats
A major drawback of likelihood-free inference methods like Dingo-IS is that if the training settings need to be changed, such as switching the waveform approximant, a new neural network must be trained using simulations generated with the updated settings. Training is not only the most time-consuming but also the most energy-intensive process of using NPE, as discussed in Sec. IV.6.
To accelerate training and improve convergence of Dingo-IS, we made a few simplifying assumptions. We set a low-frequency limit of $f_{\mathrm{min}}=6$ Hz, even though ET is expected to be sensitive down to $\sim 2$ Hz. Reducing $f_{\mathrm{min}}$ allows for processing longer signals—up to 64 s for mergers happening at $f<f_{\mathrm{max}}=256$ Hz—but requires a finer frequency resolution of $\mathrm{d}f=1/64$ Hz. This increases the input array size by a factor of 8, making training more challenging. Methods for processing longer signals with Dingo-IS have been developed in Dax:2024mcn, where they include the use of multibanding and heterodyning to handle BNS signals up to one hour in length.
Additionally, we did not consider contamination from overlapping signals or transient non-Gaussian noise, which could complicate inference (Raymond:2024xzj). Addressing these limitations in future work will help establish NPE as a robust method for processing the full range of detectable sources with ET.
A comprehensive comparison of full PE performance across different ET designs is planned for future work, including an investigation of the potential benefits of the 2L configuration, where the increased baseline between detectors may help reduce multimodalities in sky localization (Fairhurst:2009tc; Fairhurst:2017mvj; Grover:2013sha; Berry:2014jja).